OpenStudy (anonymous):

can someone help me solve -10cosx-3sinx+9=0?

OpenStudy (anonymous):

Do you want to verify this equation or simplify?? Please clarify it.

OpenStudy (anonymous):

10cos x+3 sinx =9 let 10=rsin a 3=r cos a r sin(a+x)=9 a+x=sin^{-1}(9/r) x=sin^{-1}(9/r)-a r=sqrt(109) and a= tan^{-1}(10/3)